Ackert Company's last dividend was $4.00. The dividend growth rate is expected to be constant at 1.5% for 3 years, after which dividends are expected to grow at a rate of 8.0% forever. The firm's required return (r) is 10.5%. What is the best estimate of the current stock price?

$144.07

$138.77

$146.74

$140.87
Answer #1

D1=(4*1.015)=4.06

D2=(4.06*1.015)=4.1209

D3=(4.1209*1.015)=4.1827135

Value after year 3=(D3*Growth Rate)/(Required rate-Growth Rate)

=(4.1827135*1.08)/(0.105-0.08)

=$180.6932232

Hence current price=Future dividend and value*Present value of discounting factor(rate%,time period)

=4.06/1.105+4.1209/1.105^2+4.1827135/1.105^3+$180.6932232/1.105^3

=$144.07(Approx).